Summary Report – November 05th


Summary Title: Collecting Gypsum & Seeding Mars
Author’s name: Guy Murphy
Mission Status: Active.
Sol Activity Summary: A new phase of EVA research commenced today with Andrew Wheeler locating gypsum deposits and collecting samples with Dianne. Seeds were planted in the Greenhab for the first time this season. Three cooked meals were prepared. Salmon was fish of the day. Sandy cooked the first loaf of bread in the Hab bread machine.
Look Ahead Plan: The remaining EVA’s this week will focus on completing micrometeorite collection. Andrew Wheeler will process the gypsum samples collected today in the science dome and Dianne will continue with her food waste study.
Anomalies in work: Micrometeorite study was compromised by at least one bipedal life form intruding within the study area and compacting the undisturbed study surfaces.
Weather: Another clear sunny day on Mars.
Crew Physical Status: Crew continue to be in excellent health.
EVA: See EVA report.
Reports to be filed: Sol Summary, EVA Report, Journalists Report, Science Report, Operations Reports, Greenhab Report, Photos.
Support Requested: None at present.

Operations Report – November 05th

Crew 214 Operations Report 5-11-2019
SOL: 9
Name of person filing report: Andrew Wheeler
Non-nominal systems: Opportunity, Suit 10
Notes on non-nominal systems: Opportunity brakes do not work, Suit 10 is not functioning
Generator: run
Hours run: 11h 15m
From what time last night: 1840
To what time this morning: 0655
List any additional daytime hours when the generator was run: N/A
Solar— SOC % (Before generator is run at night): 65%
Diesel Reading – 60%
Station Propane Reading – 81%
Ethanol Free Gasoline: N/A
Water (loft tank): 47 gallons
Water Meter: 1455759 units
Water (static tank): 223 gallons
Static to Loft Pump used – yes
Water in Green Hab: 225 gallons
Water in Science Dome: 0 gallons
Toilet tank emptied: no
Deimos rover used: not at hab
Hours: –
Beginning charge: –
Ending charge: –
Currently charging: –
Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: –
Beginning charge: –
Ending charge: –
Currently charging: –
Spirit rover used: yes
Hours: 108.0
Beginning charge: (Before EVA): 100%
Ending charge: (On return from EVA, before recharging): 58%
Currently charging: yes
Opportunity rover used: no
Hours: 65.5
Beginning charge: 100%
Ending charge: 100%
Currently charging: yes
Curiosity rover used: yes
Hours: 113.9
Beginning charge: (Before EVA) : 100%
Ending charge: (On return from EVA, before charging): 56%
Currently charging: yes
Notes on rovers: no
ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): no
Reason for use: N/A
Oil Added? no
ATV Fuel Used: 0 Gals
# Hours the ATVs were used today: 0
Notes on ATVs: Assigned to the director
HabCar used and why, where? N/A
CrewCar used and why, where? N/A
General notes and comments: N/A
Summary of internet: Functional
Summary of suits and radios: 5 Radios and 3 earpieces that need servicing have been identified and set aside, the battery for suit 10 is fully charged and it is Suit 10 that is not functional, further information on Suit 5,s helmet has been forwarded to Capcom.
Summary of Hab operations: All systems nominal
Summary of GreenHab operations: Atmospheric equalization to the green hab was carried out during the day, plants were watered..
Summary of ScienceDome operations: All systems nominal. Sample processing following EVA was carried out.
Summary of RAM operations: Not used
Summary of any observatory issues: N/A
Summary of health and safety issues: All crew healthy
Questions, concerns and requests to Mission Support: Between 20191104:1140 and 20191105: 1100, the micrometeorite grid has been visited and trampled by at least one smallish adult (wearing boots). This has compromised a number of the plots in the unsampled area and will affect the integrity of the experiment. Whilst this is disappointing, we realize we are on public land and, in retrospect, we should have erected a temporary sign and barrier stating “please keep clear – monitoring in progress” to prevent this. We believe future rotations may benefit from this advice. We note that our daily reports identify the specific location where our field research is undertaken.

Greenhab Report – November 05th

Crew 214 GreenHab Report 05-11-2019
GreenHab Officer: Guy Murphy
Environmental control: Door left open for ventilation during the afternoon.
Average temperatures: –
Hours of supplemental light: Automatic system
Daily water usage for crops: Watering undertaken by outpost crew.
Daily water usage for research and/or other purposes:
Water in Blue Tank – 2/3rds full
Time(s) of watering for crops: late morning.
Changes to crops: N/A
Narrative: The Greenhab was activated yesterday with the first plantings of the season.
Established plants introduced to Greenhab 04/11/2019
[3x] strawberry plants (Everbearing), in medium metal planter
[2x] small Aloe Vera, in medium metal planter, frost damaged
[1x] spicy orange thyme (thymus fragrantissimus), small terracotta pot
[1x] lemon balm scented geranium (pelargonium x Melissinum), small terracotta pot
[1x] Mint Mentha – Berries & Cream, small terracotta pot
[1x] Pineapple Mint – mentha suaveolens variegated, small terracotta pot
[1x] Ice plant – Delosperma ‘jewel of the desert garnet’, small terracotta pot
[1x] Sempervivum tectorum assorted Hen & Chicks, small terracotta pot
[1x] Rosemary ‘Barbecue’ (Rosmarinus officinalis ‘Barbecue’), large plastic pot
[1x] Rosemary ‘Spice Island’ Rosmarinus officinalis ‘Spice Island’, large plastic pot
[1x] mint, grapefruit, large plastic pot
Seeds Planted in Greenhab 04/11/2019 (Large metal planting tub)
Greek Oregano – Origanum heracleoticum
Chives – Allium schoenoprasum
Sweet Marjoram – Origanum majorana
Lemon Balm – Melissa officinalis
Sage – Salvia Officinalis
Thai Basil – Siam Queen
German/Winter Thyme – Thymus vulgaris
Harvest: None
Support/supplies needed: N/A

Journalist Report – November 05th

Tuesday 5 Nov Sol 9
by Sandy Dance
Its Sol 9 and we are starting to think: Wow! the end of our time here is coming up (at least for Dianne and me, Guy and Andrew are continuing into the next rotation). Consequently Dianne is checking through the supplies to see if there are any food containers that are nearly empty and should be replaced.
Guy spent some time in the GreenHab helping Shannon, David and Atila water the recent plantings, and sowing herbs. It will be great to have the greenhouse going, a terrific resource for those weary of ultra low humidity and temperature, and those hankering after the colour green!
Today we started a new experiment: Dianne and Andrew beetled off in the two rovers to ‘White Moon’ (there are some really wacky place names round here) further north from our micrometeorite grid to collect gypsum and associated soil. This is so Andrew can see how much water can be baked out of the soil, which is of potential interest on Mars, where there is known to be quantities of gypsum. This would be another resource for ‘in-situ resource utilization’ there. They returned with four kilos of soil and 9 gypsum samples (which look crystaline to the untrained eye), and will trial the baking tomorrow. Exciting stuff!
This afternoon I took out the Celestron astronomical telescope from the Science Dome to see what’s involved in using it at night. This is not to be confused with the Musk solar telescope, or the remote access telescope situated somewhat off campus. The Celestron is an equatorial mount, so it needs the mount axis pointing to the North star, not something I can do in the day, but I was able to calibrate the finderscope against the local hills. Very nice piece of equipment, looking forward to using it later.

EVA Report – November 05th

Crew 214 EVA Report 05-11-2019
EVA # 8
Author: Andrew Wheeler
Purpose of EVA: Geological Sampling
Start time: 0930
End time: 1200
Narrative: The sixth EVA for science saw a change in target with the change in investigation focus. Hydrated minerals are a potential source of water here on Mars and at White Moon (517800E 4254650N) the mineral gypsum (CaSO4.2H2O) is one such source. We collected both crystalline specimens and gypsum rich soils to demonstrate liberation of water through simple heating. Results are pending. On our return from White Moon, we called in to TW#2 at 518477E 4253767N to collect more micrometeorite samples and we sampled 3 plots. We have now completed 43% of the grid.
Destination: White Moon via Cow Dung Road to Brahe Highway intersection and Tank Wash on return to the hab.
Coordinates (use UTM NAD27 CONUS): 517800E 4254650N
Participants: Andrew Wheeler, Dianne McGrath
Road(s) and routes per MDRS Map: Cow Dung Road to Brahe Highway (White Moon)
Mode of travel: Spirit Rover, Curiosity Rover

Science Report – November 05th

Crew 214 Science Report 5 November 2019
SOL: 9
Crew 214 – Expedition Boomerang Crew
Submitted by Science Officer Andrew Wheeler
Science Operations: The sixth scientific EVA for the mission saw us travel to White Moon (517750E 4254520N) at the junction of Brahe Highway and Cow Dung Road in search of hydrated minerals. Immediately on arrival, gypsum crystals (CaSO4.2H2O) were observed protruding from the regolith and we were able to collect two soil samples and 9 rock samples. From these we intend to demonstrate the liberation of water through heating from both the crystals and the soil.
On our return to the the hab, we called into the Tank Wash #2 site and collected an additional 3 samples for the micrometeorite search. This is 43% of the planned samples. Unfortunately, someone has visited the grid, trampled through a number of plots and compromised the integrity of the experiment.
Green Hab: Access through the tunnel between the main hab and the green hab has again been opened for atmospheric equalization. The plants were watered by the Outpost crew.
Science Dome: Samples returned from the morning’s EVA are awaiting processing before undergoing experimental procedures within the science dome.
Closed Loop Food Waste Study: The day’s waste for Sol 9 is: 25 grams; consisting of 127 calories, 1g fat, <1g saturated fat, 69mg sodium, 70g carbohydrates, <1g fibre, 2g sugar, 5g protein.
EVA: Preparations for finalizing the micrometeorite sampling experiment were undertaken.
EVA Suit Maintenance: Nominal
Additional Activities: Photographs were taken of the compromised micrometeorite sampling grid.

Copyright © The Mars Society. All rights reserved. | Main Site